# Private Online Community Detection for Censored Block Models

### v1

@@ 2.3 Differential Privacy
$\frac{Q(\tilde{A}_{i,j}=\tilde{a}|A_{i,j}=a)}{Q(\tilde{A}_{i,j}=\tilde{a}|A_{i,j}=a^{\prime})}\leq e^{\epsilon}.$
We next present the notion of $\epsilon$-edge central differential privacy (CDP) [20].
Definition 3 ($\epsilon$-edge CDP).
An estimator $\hat{\bm{\sigma}}$ satisfies $\epsilon$-edge CDP for some ϵ+italic-$\epsilon\in\mathds{R}^{+}$, if for all pairs of adjacency matrices $\mathbf{A}$ and $\mathbf{A}^{\prime}$ that differ in one edge, we have
$\frac{\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A})=\bm{\sigma})}{\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A}^{\prime})=\bm{\sigma})}\leq e^{\epsilon},\,\forall\bm{\sigma},$ (7)
@@
It is worthwhile mentioning that Central DP is a weaker guarantee compared to the local DP, as it can be readily shown that the local DP guarantee implies central DP.
3 Online Community Change Detection
3.1 Problem Setup
Suppose we observe a sequence of undirected graphs $\{G_{t},t\in\mathbb{N}\}$ generated through the ${\mathsf{CBM}}(\bm{\sigma}_{t}^{*},p,\zeta)$ model, with $\mathbf{A}_{t}$ being the adjacency matrix of $G_{t}$, as detailed in Section 2.1. We consider the setting with known $p,\zeta$ and dynamic communities in which the true labeling vector $\bm{\sigma}_{t}^{*}$ can be time-varying. This is commonly seen in practical applications. For example, in social network data, the users may switch communities due to certain events [38].
We consider a special case, in which there is an abrupt change in the community structure at time $\nu$, which is called a change-point and takes a deterministic but unknown value. The problem can be formulated as the following hypothesis test:
$\begin{array}[]{lll}H_{0}:&\bm{\sigma}_{t}^{*}=\bm{\sigma}^{\text{pre}},&~{}~{}~{}t=1,2,\dots\\ H_{1}:&\bm{\sigma}_{t}^{*}=\bm{\sigma}^{\text{pre}},&~{}~{}~{}t=1,\dots,\nu-1,\\ &\bm{\sigma}_{t}^{*}=\bm{\sigma}^{\text{post}},&~{}~{}~{}t=\nu,\tau+1,\dots\\ \end{array}$ (8)
where $\bm{\sigma}^{\text{pre}}$ and $\bm{\sigma}^{\text{post}}$ represent the pre- and post-change community labels, respectively. We further assume $\bm{\sigma}^{\text{pre}}$ is known since it can usually be estimated from historical data collected under the pre-change (nominal) regime. Let $\mathbb{P}_{\nu}$ denote the probability measure on the observation sequence when the change-point is $\nu$, and let $\mathbb{E}_{\nu}$ denote the corresponding expectation. Also, $\mathbb{P}_{\infty}$ and $\mathbb{E}_{\infty}$ denote the probability and expectation operator when there is no change (i.e., $\nu=\infty$). For brevity, we write $\mathbb{P}$ and $\mathbb{E}$ as the probability and expectation operator when $\nu=1$.
We make the following assumptions on the change in community structures.
Assumption 1.
There exists a fixed integer $\Delta\in(0,n/2)$ such that the Hamming distance between the pre-change community label $\bm{\sigma}^{\text{pre}}$ and post-change label $\bm{\sigma}^{\text{post}}$ satisfies
$\operatorname{Ham}({\bm{\sigma}^{{\text{pre}}}},{\bm{\sigma}^{{\text{post}}}})\leq\Delta.$ (9)
When the total number of nodes $n$ is large and $\Delta$ is much smaller than $n/2$, Assumption 1 implies that the change in the community structure is sparse.
3.2 Performance Criteria
Our goal is to design a detection algorithm that is able to protect the user privacy as in Definition 2 while detecting the change as quickly as possible from streaming data. The detection is performed through a stopping time $\tau$ on the observation sequence at which the change is declared [29].
More specifically, we introduce two performance measures commonly used in online detection problems: detection delay and false alarm performances. The commonly adopted definition of Worst-case Average Detection Delay (WADD) is [23]
where $\{\mathcal{F}_{k},\ k\in\mathbb{N}\}$ is the filtration $\mathcal{F}_{k}=\sigma(G_{1},\ldots,G_{k})$, with $\mathcal{F}_{0}$ denoting the trivial sigma algebra.
We measure the false alarm performance of a detector (stopping time) $\tau$ in terms of its mean time to false alarm $\mathbb{E}_{\infty}\left[\tau\right]$, defined as Average Run Length (ARL). And we denote by $\mathcal{C}(\gamma)$ the set of all tests for which the ARL is at least $\gamma$, i.e.,
$\mathcal{C}(\gamma)=\left\{\tau:\;\mathbb{E}_{\infty}\left[\tau\right]\geq\gamma\right\}.$ (11)
The optimization problem of interest is
$\inf_{\tau\in\mathcal{C}(\gamma)}{\mathsf{WADD}}(\tau).$ (12)
Lorden showed that Page’s Cumulative Sum (CUSUM) test [28] solves the problem in (12) asymptotically; this result was strengthened to exact optimality in [26]. Under the private setting, we aim to design a detection method that attains both the data privatization and efficient change detection.
4 Proposed Method: Private Window-Limited CUSUM Test
@@
In this section, we introduce the proposed joint estimation and detection scheme, in which we perform sequential estimates for the unknown post-change community structure $\bm{\sigma}^{\text{post}}$ and use it to calculate the detection statistic. We first introduce the private mechanism that outputs privatized data through a conditional distribution that is conditional on raw data. The proposed method is summarized in Algorithm 1.
Algorithm 1 Private Window-Limited CUSUM Test for Online Community Change Detection
  Input: Sequence of adjacency matrices $\{{\mathbf{A}}_{t},t\in\mathbb{N}\}$, window size $w$, threshold $b$.
  Private mechanism: for each $A_{t,i,j}$, output the privatized sample
$\operatorname{Pr}(\tilde{A}_{t,i,j}=A_{t,i,j})$ $=\frac{e^{\epsilon}}{e^{\epsilon}+2},$
$\operatorname{Pr}(\tilde{A}_{t,i,j}=x)$ $=\frac{1}{e^{\epsilon}+2},\quad\forall x\neq A_{t,i,j}.$
  Initialize $S_{1}=\cdots=S_{w}=0$; $t=w+1$.
@@
     Calculate $\bm{\hat{\sigma}}_{t-1}$ via solving (16).
     Update $S_{t}=(S_{t-1})^{+}+\ell(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1})$.
@@
4.1 Privacy Mechanism
In this subsection, we present the graph Perturbation Mechanism that we adopted for private community detection. Furthermore, we present and analyze a noise-graph addition mechanism based on Warner’s randomized response [35]. The basic idea is to perturb the edges of the random graph (i.e., the adjacency matrix $\mathbf{A}$) to obtain a perturbed adjacency matrix $\tilde{\mathbf{A}}$ of $\mathbf{A}$, where each element $A_{i,j}$ is perturbed independently to satisfy $\epsilon$-edge LDP.
More specifically, when $A_{i,j}=x$ where $x\in\{0,+1,-1\}$, the mechanism outputs $\tilde{A}_{i,j}\sim Q_{x}(\cdot)$. We adopt a special case for graph perturbation mechanism where $Q_{x}(y)=c_{2}=\frac{1}{e^{\epsilon}+2}$ when $x\neq y$, and $Q_{x}(y)=c_{1}=1-2c_{2}=\frac{e^{\epsilon}}{e^{\epsilon}+2}$ when $x=y$. Note that for the original data distribution ${\mathsf{CBM}}(\bm{\sigma},p,\zeta)$, the perturbed elements in the adjacency matrix $\tilde{\mathbf{A}}$ are distributed as ${\mathsf{CBM}}(\bm{\sigma},\tilde{p},\tilde{\zeta})$ with (see Appendix E),
Also, note that the randomization mechanism is structured in such a way that $\tilde{A}_{i,j}=\tilde{A}_{j,i}$ and $\tilde{A}_{i,i}=0$.
4.2 Detection Procedure
We apply the detection procedure on the privatized data $\{\tilde{\mathbf{A}}_{t},t\in\mathbb{N}\}$. We adopt the recently proposed window-limited CUSUM test [36], with modifications tailored for the community estimation problem. At each time $t$, we use the current community structure estimate $\bm{\hat{\sigma}}_{t-1}$ obtained from most recent private samples $\{\tilde{\mathbf{A}}_{t-w},\ldots,\tilde{\mathbf{A}}_{t-1}\}$. Then we calculate the detection statistics recursively via
$S_{t}=(S_{t-1})^{+}+\ell(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}},\bm{\hat{\sigma}}_{t-1}),\,t\geq 1,$ (14)
with initial value $S_{0}=0$, and $(x)^{+}:=\max\{x,0\}$. In (14), the function $\ell$ is defined as
$\ell(\tilde{\mathbf{A}}_{t};\bm{\sigma},\bm{\sigma}^{\prime}):=\log\frac{p(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\prime})}{p(\tilde{\mathbf{A}}_{t};\bm{\sigma})},$ (15)
which calculates the log-likelihood ratio for each randomized sample $\tilde{\mathbf{A}}_{t}$ and is an essential element in the detection algorithms. It is worthwhile mentioning that, by definition, the estimate $\bm{\hat{\sigma}}_{t-1}$ is independent with the current sample $\tilde{\mathbf{A}}_{t}$. Such independence is essential for providing the theoretical guarantee for the detection procedure.
We then introduce how we obtain the estimator $\bm{\hat{\sigma}}_{t-1}$. We adopt the maximum likelihood estimator for windowed samples:
@@ 4.2 Detection Procedure
$\bm{\hat{\sigma}}_{t}=\arg\max_{\bm{\sigma}}\{\bm{\sigma}^{T}\left(\sum_{i=0}^{w-1}\tilde{\mathbf{A}}_{t-i}\right)\bm{\sigma}\}.$
Following Section 2.2, we define $\mathbf{Y}={\bm{\sigma}}{\bm{\sigma}}^{T}$ and use the Semidefinite Relaxation for solving $\bm{\sigma}$ (after rounding the solution of (16)):
$\hat{\mathbf{Y}}=\max_{\mathbf{Y}}$ $\ \text{tr}(\left(\sum_{i=0}^{w-1}\tilde{\mathbf{A}}_{t-i}\right)\mathbf{Y})$ (16)
s.t. $\ \mathbf{Y}\succcurlyeq\mathbf{0},Y_{i,i}=1,\forall i\in[n].$
We note that the above optimization problem can be viewed as the approximate maximum likelihood solution to a multi-view stochastic block model [39]. Moreover, it is worthwhile mentioning that we allow great flexibility in the choice of window size $w$ here. When choosing $w=1$, we use the randomized sample $\tilde{\mathbf{A}}_{t-1}$ to estimate the current community structure and then accumulate the estimated log-likelihood ratio on sample $\tilde{\mathbf{A}}_{t}$. We leave more involved discussion on the impact of window size for future work.
$T:=\inf\{t:S_{t}\geq b\},$ (17)
where $b$ is a pre-specified detection threshold chosen to satisfy the desired ARL constraint (i.e., such that $T\in\mathcal{C}(\gamma)$). We will discuss the choice of the detection threshold $b$ in Section 5.
5 Theoretical Analysis
In this section, we present the theoretical analysis for the proposed private detection procedure. We first show that the procedure is indeed $\epsilon$-edge LDP (Lemma 5.1). Then, we present theoretical guarantees for the average run length, and the worst-case average detection delay. We also compare with the information-theoretic lower bound on the detection delay in the non-private scenario (Theorem 5.1) to demonstrate the tradeoff between privacy constraints and the detection performance.
By the construction of the private mechanism, it is obvious we have the following guarantee on LDP.
Lemma 5.1 (Privacy guarantee).
The perturbation used in Section 4.1 satisfies $\epsilon$-edge LDP.
Remark 5.1.
Lemma 5.1 offers a privacy guarantee for each individual release. To evaluate the overall privacy leakage when the detection algorithm is applied to $w$ adjacency matrices, we utilize the composition property of DP. Notably, advanced composition techniques based on Rényi divergence can be employed to optimize the total privacy leakage. Hence, our analysis concludes that the cumulative privacy leakage for releasing $w$ adjacency matrices per user is bounded by $\mathcal{O}(\sqrt{w}\epsilon)$ in the context of edge DP. This paper, however, is primarily concerned with exploring the privacy ramifications for each single adjacency matrix release, rather than the cumulative effect over multiple releases.
We first present the well-known information-theoretic lower bound on the detection delay for online change detection on the above privatized data. Such lower bound will serve as the baseline, and we will compare the performance of our proposed method with such lower bound both theoretically and numerically. We denote the KL divergence between the post- and the pre-change privatized distributions as
$\tilde{I}_{0}:=\tilde{I}(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}}):=\mathbb{E}[\ell(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})],$ (18)
where ${\mathbb{E}}$ is the expectation under the post-change regime as defined above, and $\tilde{\mathbf{A}}$ is the privatized data. The known information-theoretic lower bound on the detection delay on the privatized data is a function of the KL divergence as shown below.
Theorem 5.1 (Information-Theoretic Lower Bound on WADD [32]).
A lower bound on the WADD of any test, applied on the privatized sequence of adjacency matrices $\{\tilde{\mathbf{A}}_{t},t\in\mathbb{N}\}$, that satisfies the average run length constraint $\mathbb{E}_{\infty}\left[\tau\right]\geq\gamma$ is
${\mathsf{WADD}}(\tau)\geq\frac{\log\gamma}{\tilde{I}_{0}}(1+o(1)).$
Remark 5.2.
Here we note that for given distribution parameter $p,\zeta$, the KL divergence $I(\bm{\sigma}^{\text{pre}};\bm{\sigma}^{\text{post}})$ between the post-change ${\mathsf{CBM}}(\bm{\sigma}^{\text{post}},p,\zeta)$ and pre-change ${\mathsf{CBM}}(\bm{\sigma}^{\text{pre}},p,\zeta)$ equals to
$\left(\frac{1}{2}\log\frac{1-\zeta}{\zeta}\right)p(1-2\zeta)\left[{n\choose 2}-C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}\right].$
where $C_{\bm{\sigma}^{\text{pre}},\bm{\sigma}^{\text{post}}}=\sum_{i<j}\bm{\sigma}_{i}^{\text{pre}}\bm{\sigma}_{j}^{\text{pre}}\bm{\sigma}_{i}^{\text{post}}\bm{\sigma}_{j}^{\text{post}}$. See Appendix A for detailed derivations.
5.1 Performance Analysis for the Proposed Method

### v2

@@ II-B Differential Privacy for Graphs
${Q(\tilde{A}_{i,j}=\tilde{a}|A_{i,j}=a)}\leq e^{\epsilon}{Q(\tilde{A}_{i,j}=\tilde{a}|A_{i,j}=a^{\prime})}.$
Conversely, in the central setting, privacy protections are applied to the aggregated data, focusing on preserving the presence or absence of single edges. We present the notion of $(\epsilon,\delta)$-edge central DP (CDP) [24] below. Here, $\epsilon$ is coined as privacy budget and $\delta$ represents failure probability. It is worth noting that smaller $\epsilon$, $\delta$ correspond to stronger privacy guarantee and vice versa. The setting when $\delta=0$ is referred to as pure $\epsilon$-edge CDP.
Definition 2 ($(\epsilon,\delta)$-edge CDP).
[…]
${\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A})\in\mathcal{S}})\leq e^{\epsilon}{\operatorname{Pr}(\hat{\bm{\sigma}}(\mathbf{A}^{\prime})\in\mathcal{S}})+\delta,$
@@
III Proposed Method: Privatized Adaptive
CUSUM Test
@@ III Proposed Method: Privatized Adaptive
In this section, we introduce the proposed joint estimation and detection scheme, in which we perform sequential estimates for the unknown post-change community structure $\bm{\sigma}^{\text{post}}$ and use it to calculate the detection statistic. We present two types of solution schemes for the LDP and CDP settings, respectively. Under the LDP setting, we first introduce the graph perturbation mechanism that outputs privatized data through a conditional distribution, and then introduce the Adaptive CUSUM type detection statistics utilizing the estimates of unknown post-change community $\bm{\sigma}^{\text{post}}$. Under the CDP setting, we first compute private estimates of $\bm{\sigma}^{\text{post}}$ using the original (non-private) data, and then calculate the detection statistics with added Laplace noise.
To obviate redundancy, we shall describe the first mechanism, graph perturbation mechanism, in full detail while treating the subsequent mechanism with conciseness. Our emphasis will be directed towards elucidating the main algorithmic distinctions, and their respective recovery conditions will be deliberated upon in subsequent sections. We summarize our theoretical results in Table II, which shows the tradeoffs between $(a,\zeta)$, $(\epsilon,\delta)$ as well as the computational complexity of the mechanisms.
TABLE II: Summary of the recovery threshold(s), detection delay(s), complexity, and privacy.
Mechanism Graph Perturbation Perturbation Stability
$a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}\geq$ Theorem IV.1 Theorem IV.2
Detection Delay Theorem IV.5 Theorem IV.6
Time complexity $\mathcal{O}(\operatorname{poly}(n))$ $n^{(\mathcal{O}(\log{(n)}))}$
$\epsilon$ $\Omega(\log(n))$ $\mathcal{O}(1)$
$\delta$ 0 $1/n^{2}$
III-A Detection Procedure under the LDP Setting
Privacy Mechanism (1): Graph Perturbation Mechanism
Figure 2: Illustration of the graph perturbation-based mechanism: The original edges are first randomized via a randomized response scheme, and the resulting noisy adjacency matrix is then supplied to the community recovery algorithm.
[…]
Also, the randomization mechanism is structured in such a way that $\tilde{A}_{i,j}=\tilde{A}_{j,i}$ and $\tilde{A}_{i,i}=0$. It is obvious that the above construction satisfies $\epsilon$-edge LDP in Definition 1.
Detection Procedure
We then apply the detection procedure on the privatized data $\{\tilde{\mathbf{A}}_{t},t\in\mathbb{N}\}$. We tailor the framework of the Adaptive CUSUM test [39] to our community detection problem. At each time $t$, we use the latest community structure estimate $\bm{\hat{\sigma}}_{t-1}$ obtained from the recent private sample $\tilde{\mathbf{A}}_{t-1}$ [^2] We note that it is straightforward to replace the estimator $\bm{\hat{\sigma}}_{t-1}$ as the one obtained from the most recent $w>1$ private samples $\{\tilde{\mathbf{A}}_{t-w},\ldots,\tilde{\mathbf{A}}_{t-1}\}$, which coincides with the approximate maximum likelihood solution to a multi-view stochastic block model [8, 42].. We adopt the ML estimator obtained from semidefinite relaxation as in (3). Then, we calculate the detection statistics recursively via
$S_{t}=(S_{t-1})^{+}+\log\frac{\Pr(\tilde{\mathbf{A}}_{t};\bm{\hat{\sigma}}_{t-1})}{\Pr(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}})},t\geq 1;\qquad S_{0}=0,$ (7)
where $(x)^{+}:=\max\{x,0\}$ and $\Pr(\cdot;\bm{\sigma})$ denotes the probability mass function of ${\mathsf{CBM}}(\bm{\sigma},p,\zeta)$. Here we assume the distribution parameters $p$ and $\zeta$ are known. It is worthwhile mentioning that, by definition, the estimate $\bm{\hat{\sigma}}_{t-1}$ is independent with the current sample $\tilde{\mathbf{A}}_{t}$. Such independence is essential for subsequent theoretical analysis.
$T_{\rm L}:=\inf\{t:S_{t}\geq b\},$ (8)
where $b$ is a pre-specified detection threshold chosen to meet the desired false alarm constraint. We will discuss the choice of the threshold $b$ in Section IV. The complete detection procedure is summarized in Algorithm 1.
Algorithm 1 Private Adaptive CUSUM via Graph Perturbation Mechanism
Input: Data sequence $\{{\mathbf{A}}_{t},t\in\mathbb{N}\}$, detection threshold $b$.
 Initialize $S_{1}=0$; $t=2$.
@@ III-A Detection Procedure under the LDP Setting
  Calculate $\bm{\hat{\sigma}}_{t-1}$ via solving (3) on perturbed graph $\tilde{\mathbf{A}}_{t-1}$.
  Update $S_{t}=(S_{t-1})^{+}+\log\frac{\Pr(\tilde{\mathbf{A}}_{t};\bm{\hat{\sigma}}_{t-1})}{\Pr(\tilde{\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}})}$.
@@
Remark III.1 (Unknown Parameters).
In the detection procedure, for simplicity, we make the assumption that $\bm{\sigma}^{\text{pre}}$, $p$, and $\zeta$ are known. Our detection procedure can be extended to the case with unknown parameters as well (see Appendix F for more details). In short, when these parameters are unknown, usually there is a sufficient amount of historical data sampled from the pre-change nominal scenario which can be used to estimate them to a desired accuracy. Furthermore, in the more general case where the parameters $p$ and $\zeta$ may also change, and their post-change values are unknown, we can use the most recent samples to jointly estimate $p$ and $\zeta$ and then substitute to yield the detection statistics.
III-B Detection Procedure under the CDP Setting
Privacy Mechanism (2): Perturbation Stability Mechanism
We present a stability-based community estimator $\hat{\bm{\sigma}}$ that satisfies the CDP conditions, which will then be used in the calculation of detection statistics. Given a graph represented by the adjacency matrix $\mathbf{A}$}, we first compute privately the stability of the ML estimator with respect to $\mathbf{A}$}} denoted by $d({{\mathbf{A}}})$, which is defined by the minimum number of edge modifications on $\mathbf{A}$}}, so that the estimator output on the modified graph $\mathbf{A}$$^{\prime}$ differs from that on $\mathbf{A}$}}, i.e., $\hat{\bm{\sigma}}({{\mathbf{A}}})\neq\hat{\bm{\sigma}}({{\mathbf{A}}}^{\prime})$. The formal definition of stability is given next.
Definition 3 (Stability of $\hat{\bm{\sigma}}$).
The stability of an estimator $\hat{\bm{\sigma}}$ with respect to a graph $\mathbf{A}$}} is defined as follows,
$d({\mathbf{A}})=\{\min k:\exists{\mathbf{A}}^{\prime},\text{dist}({\mathbf{A}},{\mathbf{A}}^{\prime})\leq k+1,\hat{\bm{\sigma}}({\mathbf{A}})\neq\hat{\bm{\sigma}}({\mathbf{A}}^{\prime})\},$
where $\text{dist}({\mathbf{A}},{{\mathbf{A}}}^{\prime})$ is defined as the number of differing entries between the two adjacency matrices.
If the graph $\mathbf{A}$}} is stable enough w.r.t. the recovery algorithm (i.e., if $d({\mathbf{A}})$ is larger than a threshold, which depends on $(\epsilon,\delta)$), then we release the non-private estimate $\hat{\bm{\sigma}}({{\mathbf{A}}})$; otherwise we release a random label vector. The key intuition is that from the output of a stable estimator, one cannot precisely infer the presence or absence of a single edge (thereby providing edge DP guarantee). The stability-based recovery mechanism is presented in Algorithm 2. The proposed mechanism can be readily shown to satisfy $(\epsilon,\delta)$-edge CDP in Definition 2 [14].
Algorithm 2 Perturbation Stability-based Community Recovery Algorithm
1:Input: The graph G(V,E)G({V},{E}) with adjacency matrix $\mathbf{A}$}.
2:Output: A private labelling vector $\hat{\bm{\sigma}}$.
[…]
5:if $\tilde{d}>\frac{\log{1/\delta}}{\epsilon}$ then
6:  Output $\hat{\bm{\sigma}}({{\mathbf{A}}})$
7:else
8:  Output $\perp$ (a random label vector)
9:end if
Remark III.2 (Efficient Computation).
The complexity of the stability based mechanism using semidefinite relaxation of ML estimator in (3) is $\Theta(n^{\mathcal{O}(\log(n))})$. Demonstrably, the algorithm maintains its efficacy if we opt for $\min\{{d({{\mathbf{A}}}),\mathcal{O}(\log{n})}\}$ in place of $d({{\mathbf{A}}})$. To achieve this, it is sufficient to compute $\hat{\bm{\sigma}}({{\mathbf{A}}}^{\prime})$ solely for those graphs $\mathbf{A}$$^{\prime}$ where $\text{dist}({{\mathbf{A}}},{{\mathbf{A}}}^{\prime})=\mathcal{O}(\log{n})$. Computing the distance to instability still needs to be computationally efficient. We remedy this problem by designing a computationally efficient private estimator for the distance to instability that runs in $\mathcal{O}(\log^{3}(n)\cdot\text{poly}(n)/\epsilon^{2})$, with a negligible privacy leakage term $\delta$ on the leakage as we show in the following.
From Concentration to Stability
The complexity of proving the stability of the semidefinite relaxation for ML surpasses that of the vanilla ML due to certain factors related to the optimization problem of SDP. In SDP, the ground truth label is considered as the optimal solution only under some conditions, making it challenging to establish that a solution is not a minimum bisection. Consequently, we employ the concept of concentration to establish stability. We demonstrate that all graphs within a distance of $\mathcal{O}(\log{n})$ from $G$ also exhibit concentration. Lastly, we derive a lower bound on $a\big(\sqrt{1-\zeta}-\sqrt{\zeta}\big)^{2}$ that satisfies two conditions: (1) it is sufficient for concentration and (2) it preserves concentration even when we flip up to $\Omega(\log{n})$ connections.
A Computationally Efficient Subsampling Stability Mechanism
In order to calculate the stability of our semidefinite relaxation of ML estimator in (3) efficiently, we present an algorithm by deriving a lower bound on the distance to instability $d$ as a proxy for stability, termed as “subsampling stability-based estimator”, see Fig. 3 for an illustration.
Figure 3: Illustration of the subsampling stability-based mechanism: From the original graph we generate $m$ correlated subgraphs by including each edge independently with probability $q_{s}$. A non-private community estimator is then applied to every subgraph, producing label vectors $\{\hat{\bm{\sigma}}({\mathbf{A}}_{k})\}_{k=1}^{m}$. These labelings are tallied in a histogram, and the modal labeling— provably stable under appropriately chosen $(q_{s},m,\epsilon,\delta)$—is returned as the private output.
The key idea in this algorithm is to create $m$ correlated subgraphs $\{{{\mathbf{A}}}_{1},{{\mathbf{A}}}_{2},\ldots,{{\mathbf{A}}}_{m}\}$ of the original graph $\mathbf{A}$}} where each subgraph ${{\mathbf{A}}}_{k}$ is generated by randomly subsampling with replacement of the edges in $\mathbf{A}$}} with probability $q_{s}$. We then apply our non-private ML-based estimator (3) on each subgraph ${{\mathbf{A}}}_{k}$. The labeling vectors $\hat{\bm{\sigma}}({{\mathbf{A}}}_{k})$ are then represented on a histogram. Now, define $\text{count}{(\bm{\sigma})}\triangleq|\{k\in[m]:\hat{\bm{\sigma}}({{\mathbf{A}}}_{k})=\bm{\sigma}\}|$. As shown in [16], the stability of the histogram is proportional to the difference between the most frequent bin (i.e., the mode) and the second most frequent bin. In other words, the most frequent outcome of the histogram agrees with the outcome of the original graph with high probability under an appropriate choice of the CBM parameters $(a,\zeta)$, the privacy budget $(\epsilon,\delta)$, the edge sampling probability $q_{s}$, and the number of subsampled weighted graphs $m$. We summarize the mechanism in Algorithm 3.
Algorithm 3 Subsampling-based Efficient Private Community Recovery
1:Input: The graph G(V,E)G({V},{E}) with adjacency matrix $\mathbf{A}$}.
2:Output: A private labelling vector $\hat{\bm{\sigma}}$.
[…]
4: Subsample $m$ subgraphs ${{\mathbf{A}}}_{1},{{\mathbf{A}}}_{2},\cdots,{{\mathbf{A}}}_{m}$.
5: Compute $\bar{\bm{\sigma}}=(\hat{\bm{\sigma}}({{\mathbf{A}}}_{1}),\hat{\bm{\sigma}}({{\mathbf{A}}}_{2}),\cdots,\hat{\bm{\sigma}}({{\mathbf{A}}}_{m}))$.
[…]
8:if $\tilde{d}>\frac{\log{1/\delta}}{\epsilon}$ then
9:  Output $\hat{\bm{\sigma}}({{\mathbf{A}}})=\text{mode}(\bar{\bm{\sigma}})$
10:else
11:  Output $\perp$ (a random label vector)
12:end if
Detection Procedure
Given the stability-based community estimator $\hat{\bm{\sigma}}_{t-1}$ obtained from the last sample ${\mathbf{A}}_{t-1}$, we calculate the privatized Adaptive CUSUM statistics $\tilde{S}_{t}$ as follows:
@@ A Computationally Efficient Subsampling Stability Mechanism
$\mathcal{S}_{t}=(\mathcal{S}_{t-1})^{+}+\log\frac{\Pr({\mathbf{A}}_{t};\bm{\hat{\sigma}}_{t-1})}{\Pr({\mathbf{A}}_{t};\bm{\sigma}^{\text{pre}})},\qquad\tilde{S}_{t}=\mathcal{S}_{t}+\rm{Lap}\left(\frac{4C}{\epsilon}\right),$ (9)
where $C=2\log(({1-\zeta})/{\zeta})$ is the maximum difference $\big|\log\frac{\Pr({\mathbf{A}};\bm{\hat{\sigma}}_{t-1})}{\Pr({\mathbf{A}};\bm{\sigma}^{\text{pre}})}-\log\frac{p({\mathbf{A}}^{\prime};\bm{\hat{\sigma}}_{t-1})}{p({\mathbf{A}}^{\prime};\bm{\sigma}^{\text{pre}})}\big|$ for any pair of $({\mathbf{A}},{\mathbf{A}}^{\prime})$ that differs in one edge. Note that the statistics in (9) is different from (7) in two ways: (i) the estimator $\bm{\hat{\sigma}}_{t}$ is obtained differently due to the difference between LDP and CDP settings; (ii) the non-private data ${\mathbf{A}}_{t}$ is utilized to update the statistics, consequently we add Laplace noise to the detection statistics as a common practice to ensure individual data privacy [14]. The stopping rule is constructed similarly as
$T_{\rm C}:=\inf\{t:\tilde{S}_{t}\geq\tilde{b}\}$, (10)
where $\tilde{b}=b+\rm{Lap}(\frac{2C}{\epsilon})$ is a randomized threshold based on a pre-specified constant $b$. By the DP property of the AboveThresh algorithm [15, 14], we have the above stopping rule is $\epsilon$-edge CDP.
IV Theoretical Analysis
In this section, we present a theoretical analysis for two proposed private detection frameworks. We first present sufficient and necessary conditions to ensure the exact recovery of the unknown community labels in Section IV-A. Then we present theoretical guarantees for the average run length and detection delay in Section IV-B. All proofs are deferred to the Appendices.
IV-A Conditions for Exact Community Recovery
Sufficient Condition under Graph Perturbation Mechanism